Solubility products for a series of iodides are:
AgI Ksp = 8.3 x 10-17
PbI2 Ksp = 7.1 x 10-9
BiI3 Ksp = 8.1 x 10-19
List these three compounds in order of decreasing (from highest to lowest) molar
solubility in:
a. Water
b. 0.10 M NaI
1
Expert's answer
2020-06-08T15:16:25-0400

The molar solubility of the compound, s, represents the number of moles of the substance that will dissolve in aqueous solution at a particular temperature.

For AgI s=[Ag+] and Ksp=[Ag+][I-]=s*s=s2, then s= "\sqrt{Ksp}" ="\sqrt{8.3\cdot10^{-17} }" ="9.3\cdot10^{-9}"

for PbI2 s=[Pb2+] and Ksp=[Pb2+][I-]2=s*(2s)2=4s3, then s="\sqrt[3]{Ksp\over4}=\sqrt[3]{7.1\cdot10^{-9}\over{4}}=1.2\cdot10^{-3}" ,

for BiI3 s=[Bi3+] and Ksp=[Bi3+][I-]3=s*(3s)3=27s4, then s="\sqrt[4]{Ksp\over27}=\sqrt[4]{8.1\cdot10^{-19}\over{27}}=1.3\cdot10^{-5}" .

From these considerations we get a list in order of decreasing molar solubility in water:

PbI2"\rightarrow" BiI3"\rightarrow" AgI

For 0.10 M NaI solution [I-]=0.1mol*L-1. So, we get:

For AgI s=[Ag+] and Ksp=[Ag+][I-]= s*0.1=0.1s, then s=10"\cdot" Ksp=8.3⋅10−16

for PbI2 s=[Pb2+] and Ksp=[Pb2+][I-]2=s*(0.1)2=0.01s, then s=100"\cdot" Ksp=7.1⋅10−9=7.1⋅10−7

for BiI3 s=[Bi3+] and Ksp=[Bi3+][I-]3=s*(0.1)3=0.001s, then s=1000"\cdot" Ksp=8.1⋅10−16

From these considerations we get a list in order of decreasing molar solubility in 0.10 M NaI solution:

PbI2→AgI→BiI3


Answer:

a. a list in order of decreasing molar solubility in water: PbI2→BiI3→AgI

b. a list in order of decreasing molar solubility in 0.10 M NaI solution: PbI2→AgI→BiI3